Please help with fraction equation. 5/3u=2/9
Answer by checkley75(3666)

5U/3=2/9 NOW CROSS MULTIPLY
5U*9=3*2
45U=6
U=6/45
U=2/15
PROOF
(5*2/15)/3=2/9
(10/15)/3=2/9
9(10/15)=3*2
90/15=6
6=6